A cube of mass m and height H slides with a speed v. It strikes the obstacle of height h=H/4. Find the speed of the CM of the cube just after the collision.

Ans:-51/2v/6




Grade:9

1 Answers

Chetan Mandayam Nayakar
13 years ago

L=mvH/4=Iw,I=mH2((1/6 + 1/16 + 1/4)=23mH2/48,w=mvH/4I=12v/23H, Vcm=w*H√5/4=3√5v/23